OpenStudy (anonymous):

Solve the system of equations. 6x + 2y = 10 and -6x - 2y = 12

OpenStudy (anonymous):

try adding the two equations simultaneously

OpenStudy (anonymous):

This one is actually a bit different from the others. 6x + 2y = 10 -6x - 2y = 12 When solved using any method, BOTH x and y cancel out. 0 = 22 THis is a no solution. This means that these two lines are parallel, meaning that they never intersect meaning no solutions.

OpenStudy (anonymous):

okay, i understand that! :)